# Two-Phase Channel Estimation for RIS-Assisted THz Systems with Beam Split

### v2

@@ II-A Signal and Channel Model
where $\mathbf{A}_{\mathrm{R},m}(\tilde{\boldsymbol{\psi}})=[\mathbf{a}_{\mathrm{R},m}(\tilde{\psi}_{1}),\cdots,\mathbf{a}_{\mathrm{R},m}(\tilde{\psi}_{L})]$ and $\mathbf{A}_{\mathrm{T},m}\left(\tilde{\boldsymbol{\varphi}}\right)=\left[\mathbf{a}_{\mathrm{T},m}(\tilde{\varphi}_{1}),\cdots,\mathbf{a}_{\mathrm{T},m}(\tilde{\varphi}_{L})\right]$ denote the array response matrix (ARM) of the AoDs at BS and AoAs at RIS, respectively. $\mathbf{\Sigma}_{m}=\mathrm{Diag}\left(\alpha_{1}e^{-j2\pi\tau_{1}f_{m}},\cdots,\alpha_{L}e^{-j2\pi\tau_{L}f_{m}}\right)$ is the angle-excluded matrix. Similarly, the channel $\mathbf{h}_{k}[m]\in\mathbb{C}^{1\times N_{\mathrm{R}}}$ between RIS and UE-$k$ at the $m$-th SC can be formulated as
$\mathbf{h}_{k}[m]=\sum_{j=1}^{J_{k}}{\beta_{k,j}\mathbf{a}_{\mathrm{R},m}^{\mathsf{H}}(\tilde{\vartheta}_{k,j})e^{-j2\pi\mu_{k,j}f_{m}}},$ (8)
where $J_{k}$ represents the number of propagation paths between RIS and UE-$k$, $\beta_{k,j}$, $\mu_{k,j}$ and $\tilde{\vartheta}_{k,j}$ represent the complex gain, the time delay and the AoD at RIS for the $j$-th path, respectively. By defining $\tilde{\boldsymbol{\vartheta}}_{k}=[\tilde{\vartheta}_{k,1},\cdots,\tilde{\vartheta}_{k,J_{k}}]$, we transform $\mathbf{h}_{k}[m]$ into the VAD form as
@@
where $\boldsymbol{\beta}_{k,m}\!=\!\left[\beta_{k,1}e^{-j2\pi\mu_{k,1}f_{m}},\cdots\!,\beta_{k,J_{k}}e^{-j2\pi\mu_{k,J_{k}}f_{m}}\right]^{\mathsf{T}}$ is the angle-excluded vector for the $k$-th UE. $\mathbf{A}_{\mathrm{R},m}(\tilde{\boldsymbol{\vartheta}}_{k})=[\mathbf{a}_{\mathrm{R},m}(\tilde{\vartheta}_{k,1}),\cdots,\mathbf{a}_{\mathrm{R},m}(\tilde{\vartheta}_{k,J_{k}})]$ denotes the ARM of the AoDs at RIS. Substituting (7) and (9) into (3), the cascaded channel $\mathbf{H}_{k}^{\mathrm{cas}}[m]$ can be rewritten as
$\begin{split}&\mathbf{H}_{k}^{\mathrm{cas}}[m]\\ &=\!\mathrm{Diag}(\boldsymbol{\beta}_{k,m}^{\mathsf{T}}\mathbf{A}_{\mathrm{R},m}^{\mathsf{H}}(\tilde{\boldsymbol{\vartheta}}_{k}))\mathbf{A}_{\mathrm{R},m}(\tilde{\boldsymbol{\psi}})\mathbf{\Sigma}_{m}\mathbf{A}_{\mathrm{T},m}^{\mathsf{H}}(\tilde{\boldsymbol{\varphi}})\\ &=(\mathbf{A}_{\mathrm{R},m}^{*}(\tilde{\boldsymbol{\vartheta}}_{k})\boldsymbol{\beta}_{k,m})\!\bullet\!(\mathbf{A}_{\mathrm{R},m}(\tilde{\boldsymbol{\psi}})\mathbf{\Sigma}_{m}\mathbf{A}_{\mathrm{T},m}^{\mathsf{H}}(\tilde{\boldsymbol{\varphi}}))\\ &=(\mathbf{A}_{\mathrm{R},m}^{*}(\tilde{\boldsymbol{\vartheta}}_{k})\!\bullet\!\mathbf{A}_{\mathrm{R},m}(\tilde{\boldsymbol{\psi}}))(\boldsymbol{\beta}_{k,m}\!\otimes\!\mathbf{\Sigma}_{m})\mathbf{A}_{\mathrm{T},m}^{\mathsf{H}}(\tilde{\boldsymbol{\varphi}}).\end{split}$ (10)
@@ II-B Beam Split Effect
According to (4) and (8), THz channel varies across different SCs due to the beam split effect [^1] The beam split effect can been seen as a severe version of beam squint effect in mmWave band [27].. This phenomenon is caused by the ultra-wide bandwidth and extensive antenna arrays and reflecting elements in RIS-assisted wideband THz systems. Beams generated by the frequency-independent analog beamforming vector and then reflected by the frequency-independent RIS phase shift matrix would split into totally distinct spatial directions. Typically, antenna spacings are set to half the wavelength at the central frequency $f_{c}$, i.e., $d_{\mathrm{BS}}=d_{\mathrm{RIS}}=\lambda_{c}/2$, where $\lambda_{c}=c/f_{c}$ is the wavelength of the central frequency [28]. For simplicity, we define the variables $u\triangleq\sin\tilde{u}\in[-1,1]$ and $v\triangleq\sin\tilde{v}\in[-1,1]$ to represent the physical directions. Substituting the half-wavelength configuration into (6a) and (6b), the ARVs can be rewritten as
$\mathbf{a}_{\mathrm{T},m}(u)=\frac{1}{\sqrt{N_{\mathrm{T}}}}e^{-j\pi\eta_{m}u\mathbf{n}_{t}},$ (11a)
$\mathbf{a}_{\mathrm{R},m}(v)=\frac{1}{\sqrt{N_{\mathrm{R}}}}e^{-j\pi\eta_{m}v\mathbf{n}_{r}},$ (11b)
where $\eta_{m}\triangleq f_{m}/f_{c}$ represents the relative frequency. As shown in (11), the spatial directions can be represented by
$u_{m}=\eta_{m}u,$ (12a)
$v_{m}=\eta_{m}v.$ (12b)
In the lower frequency bands such as mmWave, the subcarrier frequencies are relatively closer, which causes slight beam squint with $\eta_{m}\approx 1$ which means that the ARVs at all SC frequencies are almost identical[28, 20]. However, due to the ultra large bandwidth in THz band, $\eta_{m}$ deviates significantly from one. Therefore, the difference between physical directions and spatial directions becomes significant and cannot be ignored.
@@
In wideband RIS-assisted THz systems, beam split effect occurs at both BS and RIS, which leads to severe performance degradation in channel estimation algorithms [27]. Therefore, it is crucial to develop an effective channel estimation scheme to acquire accurate CSI under beam split effect.
III Channel Estimation
III-A Channel Estimation Protocol
@@
Traditional channel estimation algorithms require significant redesign of the codebook, as conventional dictionaries fail to cover additional angle information introduced by beam split effect [17]. This additional angle information will complicates the design of dictionaries. Furthermore, the beam split effect disperses the channel for each subcarrier into different spatial directions, which requires well-designed channel estimation algorithms to operate on each subcarrier individually, which can significantly increase computational complexity.
@@
To address these challenges, we propose a novel low-complexity channel estimation protocol, which is divided into two main phases, as depicted in Fig. 1. The first phase, which is the most challenging aspect of the protocol, mainly focuses on estimating the full CSI for two selected subcarriers, which includes estimation of cascaded channel, the AoDs at BS, and the coupled angles at RIS. Due to its complexity, we further divide this phase into two steps. In Step 1, we employ sparse recovery algorithms to estimate the cascaded channel for the two subcarriers. In Step 2, we focus on heuristic algorithms for common angles estimation. Given the full CSI estimation from the first phase, we then employ a simple LS algorithm to estimate the channels for the remaining subcarriers.
III-B Cascaded Channel Estimation
@@
In this subsection, we focus on the estimation of cascaded channel for one specific subcarrier. In the following steps, we first present the angular domain sparse representation in RIS-assisted THz channels. Then, we introduce the simplified dictionary, i.e., CBS dictionary. Based on this simplified dictionary, we reformulate the channel estimation problem as a sparse recovery problem. Finally, we provide a detailed analysis of existing sparse recovery algorithms.
III-B1 Angular Domain Sparse Representation
@@
$\mathbf{H}_{k}^{\mathrm{cas}}[m]=\tilde{\mathbf{\Xi}}_{\mathrm{R},m}\left(\bar{\boldsymbol{\beta}}_{k,m}\otimes\bar{\mathbf{\Sigma}}_{m}\right)\mathbf{C}_{\mathrm{T},m}^{\mathsf{H}},$ (13)
where $\tilde{\mathbf{\Xi}}_{\mathrm{R},m}\triangleq\mathbf{C}_{\mathrm{R},m}^{*}\bullet\mathbf{C}_{\mathrm{R},m}\in\mathbb{C}^{N_{\mathrm{R}}\times Q_{\mathrm{R}}^{2}}$ is the coupled dictionary, $\bar{\mathbf{\Sigma}}_{m}\in\mathbb{C}^{Q_{\mathrm{R}}\times Q_{\mathrm{T}}}$ and $\bar{\boldsymbol{\beta}}_{k,m}\in\mathbb{C}^{Q_{\mathrm{R}}}$ are the sparse formulation with $L$ and $J_{k}$ none zero entries corresponding to the angle-excluded matrix of BS-RIS channel and the angle-excluded vector of RIS-UE channel, respectively. Additionally, the overcomplete dictionaries of the BS and RIS directions can be expressed as
@@
III-B2 Simplified Dictionary
Note that the coupled dictionary $\tilde{\mathbf{\Xi}}_{\mathrm{R},m}$ in (13) is derived from the transposed Khatri-Rao product, which contains a significant number of redundant columns due to the fixed sampling intervals. For narrowband channel model, this will result in only $Q_{\mathrm{R}}$ unique columns which are exactly the first $Q_{\mathrm{R}}$ columns [9]. However, for wideband THz channel with beam split effect, this simplification is no longer suitable due to the frequency-dependent nature of the ARVs.
@@ III-B2 Simplified Dictionary
The number of unique columns of dictionary $\tilde{\mathbf{\Xi}}_{\mathrm{R},m}$ is $Q_{B}=2Q_{R}-1$, and the coupled dictionary can be simplified by including only the first and the last $Q_{\mathrm{R}}$ columns of the coupled dictionary $\tilde{\mathbf{\Xi}}_{\mathrm{R},m}$, which can be given as
@@
where $\mathbf{\Xi}_{\mathrm{R},m}\in\mathbb{C}^{N_{\mathrm{R}}\times Q_{\mathrm{B}}}$ is the simplified dictionary.
Figure 3: Coupled dictionary simplification diagram.
@@
We first define the parameter sequence $\bar{\vartheta}_{q_{1}}=\delta\cdot q_{1}-(Q+1)/{Q},\forall q_{1}=1,\dots,Q_{\mathrm{R}}$ which corresponds one-to-one with $Q_{\mathrm{R}}$ ARV columns in the dictionary $\mathbf{C}_{\mathrm{R},m}$. Similarly, we define the sequence $\delta_{q_{2}}=-\delta\cdot q_{2}+(Q+1)/Q$ to correspond to $Q_{R}$ ARV columns in the conjugate dictionary $\mathbf{C}_{\mathrm{R},m}^{*}$. Thus, the column vectors of $\mathbf{C}_{\mathrm{R},m}^{*}\bullet\mathbf{C}_{\mathrm{R},m}$ can be expressed as $\mathbf{a}_{\mathrm{R},m}(\bar{\vartheta}_{q_{1}}-\bar{\vartheta}_{q_{2}})$ for $q_{1},q_{2}=1,\dots,Q_{\mathrm{R}}$. The $j$-th column of the coupled dictionary $\tilde{\boldsymbol{\xi}}_{\mathrm{R},m,j}$ can be expressed as
where $j=\left(q_{2}-1\right)Q_{\mathrm{R}}+q_{1},\forall q_{1},q_{2}=1,\dots,Q_{\mathrm{R}}$. Based on the fact that $q_{1}-q_{2}\in\left[1-Q_{\mathrm{R}},Q_{\mathrm{R}}-1\right]$, the number of unique columns of coupled dictionary $\tilde{\mathbf{\Xi}}_{\mathrm{R},m}$ is $Q_{\mathrm{B}}=2Q_{\mathrm{R}}-1$.
Then, we simplify the coupled dictionary $\tilde{\mathbf{\Xi}}_{\mathrm{R},m}$ step by step from (a) to (d) as shown in Fig. 3. As depicted in step (a), we reshape the overcomplete dictionary $\tilde{\mathbf{\Xi}}_{\mathrm{R},m}$ of length $Q_{\mathrm{R}}^{2}$ ARV columns into a three-dimensional matrix where each column represents an ARV. Given that ARV vectors can be uniquely determined by known angles in (16), we use the sequences $q_{1},q_{2}=1,\dots,Q_{\mathrm{R}}$ to index the corresponding ARV columns. In step (b), we form the index matrix $\mathbf{Q}\in\mathbb{N}^{Q_{\mathrm{R}}\times Q_{\mathrm{R}}}$ with elements $[\mathbf{Q}]_{i,j}=q_{1}(i)-q_{2}(j)$ for $i,j=1,\dots,Q_{\mathrm{R}}$. It is noticed that due to the cyclic shift property of this dictionary, the dictionary base can be represented entirely by combining just the first and last rows. This means that we can reconstruct the original dictionary $\tilde{\mathbf{\Xi}}_{\mathrm{R},m}$ with a sliding window of length $Q_{\mathrm{R}}$. Therefore, the simplified CBS dictionary $\mathbf{\Xi}_{\mathrm{R},m}$ can be given as in (15). Moreover, substituting $q_{2}=Q_{\mathrm{R}}$ into (16), the $q$-th column where $q=1,\dots Q_{\mathrm{B}}$ of $\mathbf{\Xi}_{\mathrm{R},m}$ can be formulated as
III-B3 Sparse Recovery Reformulation

### v3

@@ II-A Signal and Channel Model
where $\mathbf{A}_{\mathrm{R},m}(\tilde{\boldsymbol{\psi}})=[\mathbf{a}_{\mathrm{R},m}(\tilde{\psi}_{1}),\cdots,\mathbf{a}_{\mathrm{R},m}(\tilde{\psi}_{L})]$ and $\mathbf{A}_{\mathrm{T},m}\left(\tilde{\boldsymbol{\varphi}}\right)=\left[\mathbf{a}_{\mathrm{T},m}(\tilde{\varphi}_{1}),\cdots,\mathbf{a}_{\mathrm{T},m}(\tilde{\varphi}_{L})\right]$ denote the array response matrix (ARM) of the DoDs at BS and DoAs at RIS, respectively. $\mathbf{\Sigma}_{m}=\mathrm{Diag}\left(\alpha_{1}e^{-\jmath 2\pi\tau_{1}f_{m}},\cdots,\alpha_{L}e^{-\jmath 2\pi\tau_{L}f_{m}}\right)$ is the angle-excluded matrix.
Similarly, the channel $\mathbf{h}_{k}[m]\in\mathbb{C}^{1\times N_{\mathrm{R}}}$ between RIS and $k$-th UE at the $m$-th SC can be formulated as
$\mathbf{h}_{k}[m]=\sum_{j=1}^{J_{k}}{\beta_{k,j}\mathbf{a}_{\mathrm{R},m}^{\mathsf{H}}(\tilde{\vartheta}_{k,j})e^{-\jmath 2\pi\mu_{k,j}f_{m}}},$ (8)
where $J_{k}$ represents the number of propagation paths between RIS and $k$-th UE, $\beta_{k,j}$, $\mu_{k,j}$ and $\tilde{\vartheta}_{k,j}$ represent the complex gain, the time delay and the DoD at RIS for the $j$-th path, respectively. By defining $\tilde{\boldsymbol{\vartheta}}_{k}=[\tilde{\vartheta}_{k,1},\cdots,\tilde{\vartheta}_{k,J_{k}}]$, $\mathbf{h}_{k}[m]$ can be rewritten as
@@
where $\boldsymbol{\beta}_{k,m}\!=\!\left[\beta_{k,1}e^{-\jmath 2\pi\mu_{k,1}f_{m}},\cdots\!,\beta_{k,J_{k}}e^{-\jmath 2\pi\mu_{k,J_{k}}f_{m}}\right]^{\mathsf{T}}$ is the angle-excluded vector for the $k$-th UE. $\mathbf{A}_{\mathrm{R},m}(\tilde{\boldsymbol{\vartheta}}_{k})=[\mathbf{a}_{\mathrm{R},m}(\tilde{\vartheta}_{k,1}),\cdots,\mathbf{a}_{\mathrm{R},m}(\tilde{\vartheta}_{k,J_{k}})]$ denotes the ARM of the DoDs at RIS. Substituting (7) and (9) into (3), the cascaded channel $\mathbf{H}_{k}^{\mathrm{cas}}[m]$ can be rewritten as
$\mathbf{H}_{k}^{\mathrm{cas}}[m]=(\mathbf{A}_{\mathrm{R},m}^{*}(\tilde{\boldsymbol{\vartheta}}_{k})\!\bullet\!\mathbf{A}_{\mathrm{R},m}(\tilde{\boldsymbol{\psi}}))(\boldsymbol{\beta}_{k,m}\!\otimes\!\mathbf{\Sigma}_{m})\mathbf{A}_{\mathrm{T},m}^{\mathsf{H}}(\tilde{\boldsymbol{\varphi}}).$ (10)
@@ II-B Beam Split Effect
According to (4) and (8), THz channel varies across different SCs due to the beam split effect [^3] The beam split effect can been seen as a severe version of beam squint effect in mmWave band [30].. This phenomenon is caused by the ultra-wide bandwidth and extensive antenna arrays and reflecting elements in RIS-assisted wideband THz systems. The beams generated by the frequency-independent analog beamforming vector and then reflected by the frequency-independent RIS phase shift matrix would split into totally distinct spatial directions. Typically, antenna spacings are set to half the wavelength at the central frequency $f_{c}$, i.e., $d_{\mathrm{BS}}=d_{\mathrm{RIS}}=\lambda_{c}/2$, where $\lambda_{c}=c/f_{c}$ is the wavelength of the central frequency [31]. For simplicity, we define the variables $u\triangleq\sin\tilde{u}\in[-1,1]$ and $v\triangleq\sin\tilde{v}\in[-1,1]$ to represent the physical directions. Substituting the half-wavelength configuration into (6a) and (6b), the ARVs can be rewritten as
$\mathbf{a}_{\mathrm{T},m}(u)=\frac{1}{\sqrt{N_{\mathrm{T}}}}e^{-\jmath\pi\eta_{m}u\mathbf{n}_{t}},$ (11a)
$\mathbf{a}_{\mathrm{R},m}(v)=\frac{1}{\sqrt{N_{\mathrm{R}}}}e^{-\jmath\pi\eta_{m}v\mathbf{n}_{r}},$ (11b)
where $\eta_{m}\triangleq f_{m}/f_{c}$ represents the relative frequency. As shown in (11), the spatial directions can be represented by $u_{m}=\eta_{m}u$ and $v_{m}=\eta_{m}v$. In the lower frequency bands, such as mmWave, the SC frequencies are relatively closer, resulting in only minor deviations of $\eta_{m}$ from one. Consequently, beam squint remains limited, and the ARVs exhibit only slight variations across different SCs [31, 20]. In contrast, the ultra-large bandwidth in THz band leads to a significantly larger deviation in $\eta_{m}$ compared to mmWave, making the discrepancy between physical and spatial directions become pronounced and non-negligible.
@@
In wideband RIS-assisted THz systems, beam split effect occurs at both BS and RIS, resulting in severe performance degradation in channel estimation algorithms [30]. Therefore, it is crucial to develop an effective channel estimation scheme to acquire accurate CSI under beam split effect.
II-C Channel Estimation Protocol
@@
Traditional channel estimation algorithms require substantial modifications to the dictionaries, as conventional designs fail to capture the additional angular information introduced by the beam split effect [17]. Furthermore, the beam split effect disperses the channel of each SC into different spatial directions, thereby necessitating well-designed channel estimation algorithms that operate independently on each SC, which significantly increases the computational complexity.
@@
To address these challenges, we propose the following low-complexity channel estimation protocol, which is divided into two main phases, as depicted in Fig. 1. Phase I mainly focuses on estimating the full CSI for two selected SCs, which includes estimation of cascaded channels, the DoDs at BS, and the coupled angles at RIS. Specifically, we further divide this phase into two steps. In Step 1, we employ sparse recovery algorithms to estimate the cascaded channel for the selected two SCs. In Step 2, we focus on the heuristic algorithms for common angles estimation. Given the full CSI estimation from Phase I, we then employ a simple LS algorithm to estimate the channels for the remaining SCs in Phase II.
III Cascaded Channel Estimation for two SCs
@@
In this section, we focus on the estimation of the cascaded channel for two specific SCs. We begin by presenting the angular-domain sparse representation of RIS-assisted THz channels. Next, we introduce the simplified dictionary, referred to as the CBS dictionary. Based on this simplified dictionary, we reformulate the channel estimation problem as a sparse recovery problem. Finally, we provide a comprehensive analysis of existing sparse recovery algorithms.
III-A Angular Domain Sparse Representation
@@
$\mathbf{H}_{k}^{\mathrm{cas}}[m]=\tilde{\mathbf{\Xi}}_{\mathrm{R},m}\bar{\mathbf{X}}_{k,m}\mathbf{C}_{\mathrm{T},m}^{\mathsf{H}},$ (12)
where $\tilde{\mathbf{\Xi}}_{\mathrm{R},m}\triangleq\mathbf{C}_{\mathrm{R},m}^{*}\bullet\mathbf{C}_{\mathrm{R},m}\in\mathbb{C}^{N_{\mathrm{R}}\times Q_{\mathrm{R}}^{2}}$ is the coupled dictionary. Moreover, $\bar{\mathbf{X}}_{k,m}\triangleq\bar{\boldsymbol{\beta}}_{k,m}\otimes\bar{\mathbf{\Sigma}}_{m}\in\mathbb{C}^{Q_{\mathrm{R}}^{2}\times Q_{\mathrm{T}}}$, where $\bar{\mathbf{\Sigma}}_{m}\in\mathbb{C}^{Q_{\mathrm{R}}\times Q_{\mathrm{T}}}$ and $\bar{\boldsymbol{\beta}}_{k,m}\in\mathbb{C}^{Q_{\mathrm{R}}}$ are the sparse formulation with $L$ and $J_{k}$ non-zero entries corresponding to the angle-excluded matrix of BS-RIS channel and the angle-excluded vector of RIS-UE channel, respectively. Consequently, $\bar{\mathbf{X}}_{k,m}$ is an $L\times J_{k}$-sparse matrix with its non-zero entries corresponding to the set $\{\!\bar{x}_{k\!,m\!,l\!,j}|\bar{x}_{k\!,m\!,l\!,j}\!\triangleq\!\alpha_{l}\beta_{k,\!j}e^{\!-\jmath 2\pi\!\left(\!\tau_{l}\!+\!\mu_{k,j}\!\right)\!f_{m}}\!,\!l\!=\!1,\!\dots\!,\!L,\!j\!\!=\!1,\!\dots\!,\!J_{k}\!\}$. Additionally, the overcomplete dictionaries of the BS and RIS directions can be expressed as
@@
III-B Simplified Dictionary
Note that the coupled dictionary $\tilde{\mathbf{\Xi}}_{\mathrm{R},m}$ contains only $Q_{\mathrm{R}}$ unique columns which are exactly the first $Q_{\mathrm{R}}$ columns in narrowband systems[9]. However, for wideband THz channel with beam split effect, this simplification is no longer suitable due to the frequency-dependent nature of the ARVs.
@@ III-B Simplified Dictionary
The dictionary $\tilde{\mathbf{\Xi}}_{\mathrm{R},m}$ contains $Q_{B}=2Q_{R}-1$ unique columns, and the coupled dictionary can be simplified by including only the first and the last $Q_{\mathrm{R}}$ columns of the coupled dictionary $\tilde{\mathbf{\Xi}}_{\mathrm{R},m}$, which can be expressed as
@@
where $\mathbf{\Xi}_{\mathrm{R},m}\in\mathbb{C}^{N_{\mathrm{R}}\times Q_{\mathrm{B}}}$ is the simplified dictionary, and the $q$-th column where $q=1,\dots Q_{\mathrm{B}}$ of $\mathbf{\Xi}_{\mathrm{R},m}$ can be formulated as
@@
The proof is provided in Appendix A. ∎
III-C Sparse Recovery Reformulation